# SQL-Factory: A Multi-Agent Framework for High-Quality and Large-Scale SQL Generation

### v3

@@ Abstract.
High quality SQL corpus is essential for intelligent database. For example, Text-to-SQL requires SQL queries and correspond natural language questions as training samples. However, collecting such query corpus remains challenging in practice due to the high cost of manual annotation, which highlights the importance of automatic SQL generation. Despite recent advances, existing generation methods still face limitations in achieving both diversity and cost-effectiveness. Besides, many methods also treat all tables equally, which overlooks schema complexity and leads to under-utilization of structurally rich tables. To address these issues, this paper proposes a multi-agent framework for high-quality and large-scale SQL generation, dubbed SQL-Factory. It decomposes the generation process into three collaborative teams: the Generation Team explores diverse query structures using a powerful language model, the Expansion Team scales promising patterns via a lightweight language model, and the Management Team adaptively schedules the workflow and evaluates the quality of synthesized queries. This modular framework ensures a balanced trade-off between diversity, scalability, and generation cost. We apply SQL-Factory to four widely used benchmarks and generate over 300,000 SQL queries with less than \$200 API cost. Our generated queries achieve higher diversity compared to other methods, and extensive experiments demonstrate that the generated queries significantly improve the model performance in various downstream tasks.
@@ 1. Introduction
A high-quality SQL corpus is essential for advancing intelligent database technologies, such as cardinality estimation (Negi et al., 2023; Li et al., 2023b; Wang et al., 2023), configuration tuning (Zhang et al., 2019; Li et al., 2025b; Van Aken et al., 2017; Bianchi et al., 2024), natural language interfaces (Giaquinto et al., 2023; Li et al., 2024; Sun et al., 2023). For example, to train a learning-based configuration tuning system, it is necessary to observe how different SQL queries interact with various configuration settings, which requires a large number of diverse SQL workloads. Besides, since SQL can be easily translated into natural language, it also facilitates the construction of datasets for tasks like Text-to-SQL.
@@
High generation cost. Currently learning-based method, especially effective large language model (LLM)-based SQL generation methods often depend on either proprietary APIs or heavyweight open-source deployments to achieve reliable quality, leading to high costs in API usage or deployment. For instance, OmniSQL (Li et al., 2025a) demonstrates the feasibility of large-scale SQL generation by using a mixture of locally deployed models, including those with over 30 billion parameters. This approach significantly increases deployment and maintenance costs.
To address these limitations, we introduce SQL-Factory, a multi-agent framework designed for high-quality and large-scale SQL query generation. SQL-Factory leverages three collaborative agent teams: the Management Team, the Generation Team, and the Expansion Team, each responsible for a distinct aspect of the generation process. The Generation Team explores structurally diverse and expressive SQL patterns using powerful large-scale language models, ensuring sufficient variety across query forms and constructs. The Expansion Team complements this by scaling the workload through lightweight and cost-effective local models, which generate semantically coherent variants based on representative seed queries, significantly reducing reliance on expensive models. Finally, the Management Team evaluates query quality and schema-level distribution in real time, guiding scheduling decisions to ensure that structurally complex or underrepresented schema components are adequately covered. This division of responsibilities enables SQL-Factory to achieve a promising trade-off between query diversity, generation efficiency, and schema-level balance.
@@ 2.1. Data-Driven Intelligent Database
Another important line of work focuses on converting natural language to SQL (NL2SQL). A wide range of methods (Pourreza and Rafiei, 2023; Wang et al., 2025b; Gao et al., 2024; Zhang et al., 2024; Xu et al., 2025) train neural models on aligned natural language and SQL query pairs, enabling users to express complex information needs in plain language and receive executable queries as output.
@@ 2.3. LLM-based Multi-Agent System
In the context of databases, LLM-based multi-agent approaches have recently emerged as a powerful solution for tasks requiring high modularity and semantic precision. For instance, in Text-to-SQL translation, multiple agents can collaboratively handle subtasks such as schema linking, query generation, and final SQL refinement, leading to more accurate SQL outputs (Wang et al., 2025b; Xie et al., 2024). In the realm of database anomaly diagnosis, systems like D-Bot (Zhou et al., 2024) leverage large language models to automatically analyze and identify root causes of anomalies by extracting knowledge from diagnostic documents. Additionally, ROMAS (Huang et al., 2024) introduces a role-based multi-agent framework for database monitoring and planning, featuring agents assigned to specific roles (planner, monitor, and worker) collaboratively facilitate dynamic task allocation, self-monitoring, and adaptive re-planning in complex data environments. These examples collectively demonstrate the potential of multi-agent architectures to enhance reasoning modularity, improve decision accuracy, and integrate system-level feedback more effectively. Despite these advances, the application of LLM-based multi-agent systems to large-scale SQL generation remains underexplored.
@@ 3. System Overview
Generation Team. The Generation Team is designed to maximize structural diversity, tackling the problem of repetitive or overly templated query outputs. It comprises the Table Selection Agent and the Generation Agent. The Table Selection Agent performs schema-aware planning by prioritizing structurally rich or underrepresented tables using schema statistics. Then, the Generation Agent employs a powerful LLM to generate novel and semantically valid queries from scratch. These queries often include complex constructs such as nested subqueries, aggregations, and multi-table joins. By focusing on pattern exploration, the Generation Team ensures the introduction of high-diversity query structures that guide subsequent expansion.
@@ 4.1.1. Management Agent
Specifically, the system begins in the exploration phase, where the Generation Team synthesizes queries for underrepresented tables. After a fixed number of generation cycles, control shifts to the Expansion Team to exploit learned patterns. During this exploitation phase, the Management Agent continuously monitors evaluation information from Critical Agent. If query redundancy increases significantly or executability declines, it triggers a transition back to exploration. This alternating control mechanism allows SQL-Factory to incrementally grow a dataset that is both diverse and scalable, while preventing the generation process from collapsing into repetitive or overly common query patterns.
@@ 4.1.2. Critical Agent
The Critical Agent serves as the evaluation engine of SQL-Factory. Its primary function is to assess the quality of newly generated or expanded SQL queries on the fly before they are admitted into the SQL Pool. This agent operates as a shared component used by both the Generation Team and the Expansion Team, and its evaluation result also plays a crucial role in guiding the Management Agent’s scheduling decisions.
@@
For SQL queries from the Generation Team, the powerful large model ensures a degree of novelty and exploratory coverage. Therefore, we focus primarily on syntactic and semantic validity. Given a set of queries $\mathcal{Q}_{t}=\{q_{t}^{1},q_{t}^{2},\dots,q_{t}^{n}\}$ at time step $t$, the Critical Agent call the Execution Tool $\text{Tool}_{\text{E}}$ to compile each query on the target database schema and reports a binary executability flag. The overall evaluation for this batch is defined as:
@@
    (5) $\text{Sim}_{\text{Ebd}}(q,q^{\prime})=\frac{\text{Ebd}(q)\cdot\text{Ebd}(q^{\prime})}{||\text{Ebd}(q)||\ ||\text{Ebd}(q^{\prime})||},$
    where $\text{Ebd}(q)$ denotes the vector representation of query $q$ encoded by the BERT encoder, and |||||$\cdot$||| | ⋅ | | represents the norm of the vector.
By integrating the above three similarity metrics, this hybrid scoring function provides a more comprehensive assessment of query similarity. It is particularly effective in identifying subtle redundancies among expanded SQL queries, and helps ensure that only genuinely novel and diverse queries are admitted into the SQL Pool.
Finally, the evaluation for the Expansion Team is defined as follows to systematically evaluate executability and diversity.
@@ 4.2.1. Table Selection Agent
The agent operates over two sources of information: (1) the full schema of the target database, including all table names and their respective column types; and (2) a statistical summary of SQL usage derived from the SQL Pool, such as the number of queries referencing each table. These metrics are extracted and structured via a dedicated Statistical Tool $\text{Tool}_{S}$, which computes table-level query coverage.
@@ 4.2.2. Generation Agent
The Generation Agent is responsible for synthesizing new SQL queries in a schema-aware and open-ended manner. It takes as input the set of tables $\mathcal{T}^{\prime}$ selected by the Table Selection Agent and aims to produce queries that are structurally diverse. This process constitutes the core exploration step in SQL-Factory, enabling the discovery of novel query patterns that are unlikely to emerge from simple perturbations of existing SQL.
@@ 4.3. Expansion Team
The expansion process follows a two-step workflow. First, the Seed Selection Agent identifies representative queries from the SQL Pool that involve under-covered tables or patterns. Then, the Expansion Agent learns from these queries and generates structurally similar but lexically and semantically varied SQL queries that preserve executability. Since this phase focuses on local exploration around known patterns, it enables SQL-Factory to rapidly increase the number of high-quality queries while maintaining system efficiency. Together, these two agents complements the structural exploration performed by the Generation Team and allows SQL-Factory to produce a large volume of diverse and executable queries at scale with low cost.
@@ 4.3.2. Expansion Agent
The Expansion Agent generates new SQL queries by leveraging representative seed queries retrieved from the Seed Selection Agent. Unlike the Generation Agent, which synthesizes queries from scratch based on schema information, the Expansion Agent focuses on local exploration. Its goal is to generate syntactically diverse and executable queries by learning from the structure or intent of seed queries. This agent enables the system to scale up query volume without incurring the computational cost of large-scale model inference.
@@ 5. Experiments
In this section, we conduct extensive experiments to demonstrate the efficiency of the SQL-Factory framework from multiple perspectives, including diversity analysis, case studies, and its effectiveness on downstream tasks.
@@ 5.1.2. Benchmarks
To evaluate the effectiveness and generality of SQL-Factory, we select four typical database benchmarks for synthesizing queries. (1) TPC-DS (Nambiar and Poess, 2006) is an industry-standard benchmark for decision support systems, featuring a complex schema with 24 tables and various star-join patterns. It serves as a rigorous testbed for generating analytical SQL queries. (2) IMDB (Maas et al., 2011) is a movie-related database, which includes 21 tables on movies and actors. Its real-world structure and moderate size make it suitable for generating realistic join-heavy queries. (3) SPIDER (Yu et al., 2018) is a large-scale text-to-SQL dataset consisting of 200 databases with highly diverse and normalized schemas. We use the SPIDER schema collection as a target set for multi-domain query generation. (4) BIRD (Li et al., 2023a) is a recent benchmark introduced to evaluate database-instruction reasoning in realistic and complex environments. We use BIRD both as a target for SQL generation and as a downstream testbed for Text-to-SQL evaluation, particularly to assess robustness under schema ambiguity. Across all benchmarks, we parse the schema into structured representations including table names, column types, and foreign key relations. These are used as inputs for SQL-Factory’s Generation and Expansion workflows.
To evaluate the effectiveness of the generated SQL queries in improving representation learning, we adopt three widely-used query clustering benchmarks used in PreQR (Tang et al., 2022): IIT Bombay (Chandra et al., 2015), UB Exam, and PocketData (Kennedy et al., 2016). These datasets contain real-world SQL queries collected from academic and industry environments, and are commonly used to evaluate unsupervised similarity learning methods for SQL.
@@ 5.1.3. Baseline
SQL Generation Baselines. To compare the quality and diversity of generated SQL queries, we consider both rule-based and learning-based generation methods. We include SQLSmith (Seltenreich, 2016), a grammar-driven query generator that produces random SQLs over a given schema, and two representative learning-based methods: LearnedSQLGen (Zhang et al., 2022), which learns to generate queries via reinforcement learning, and OmniSQL (Li et al., 2025a), which leverages large language models to synthesize large-scale queries. Additionally, we exclude template-based methods (Nambiar and Poess, 2006; Poess and Floyd, 2000; Leis et al., 2018) from our comparison, as they rely heavily on predefined templates and are not suitable for generating massive and structurally diverse SQL workloads.
@@
Query Clustering Baselines. For the SQL clustering task, we follow the baseline setup from PreQR (Tang et al., 2022) and evaluate several established similarity models. These include statistical and heuristic approaches such as Aouiche (Aouiche et al., 2006), Aligon (Aligon et al., 2014), and Makiyama (Makiyama et al., 2015), the pretrained model-based PreQR, and two general-purpose language models: BERT-Base and BERT-Large (Devlin et al., 2019). We further include versions of BERT fine-tuned with contrastive learning on SQL-Factory data to highlight the effectiveness of our generated queries in improving SQL representation learning.
@@ 5.1.4. Implementation Details
Expansion Agent. For the Expansion Agent, we utilize Qwen2.5-Coder-14B (Hui et al., 2024), a locally deployed open-source model optimized for code and SQL-related instruction following. It is served using vLLM to support efficient batch decoding at scale, enabling rapid query expansion. It is also use a temperature of 0.8 for diversity.
@@ 5.2.3. Comparative Analysis on Query Diversity
We conduct a comparative study to assess the query diversity achieved by SQL-Factory against three representative SQL generation methods: SQLSmith (Seltenreich, 2016), a grammar-based random query generator; LearnedSQLGen (Zhang et al., 2022), which employs a reinforcement learning framework guided by cost-based feedback to generate queries through a finite-state abstraction; and OmniSQL (Li et al., 2025a), a LLM-based framework that leverages prompt engineering to generate millions of executable SQL queries.

### v4

@@ Abstract.
High quality SQL corpus is essential for intelligent database. For example, Text-to-SQL requires SQL queries and corresponding natural language questions as training samples. However, collecting such query corpus remains challenging in practice due to the high cost of manual annotation, which highlights the importance of automatic SQL generation. Despite recent advances, existing generation methods still face limitations in achieving both diversity and cost-effectiveness. Besides, many methods also treat all tables equally, which overlooks schema complexity and leads to under-utilization of structurally rich tables. To address these issues, this paper proposes a multi-agent framework for high-quality and large-scale SQL generation, dubbed SQL-Factory. It decomposes the generation process into three collaborative teams: the Generation Team explores diverse query structures using a powerful language model, the Expansion Team scales promising patterns via a lightweight language model, and the Management Team adaptively schedules the workflow and evaluates the quality of synthesized queries. This modular framework ensures a balanced trade-off between diversity, scalability, and generation cost. We apply SQL-Factory to four widely used benchmarks and generate over 300,000 SQL queries with less than \$200 API cost. Our generated queries achieve higher diversity compared to other methods, and extensive experiments demonstrate that the generated queries significantly improve the model performance in various downstream tasks.
@@ 1. Introduction
A high-quality SQL corpus is essential for advancing intelligent database technologies, such as cardinality estimation (Negi et al., 2023; Li et al., 2023b; Wang et al., 2023), configuration tuning (Zhang et al., 2019; Li et al., 2025b; Van Aken et al., 2017; Bianchi et al., 2024) and natural language interfaces (Giaquinto et al., 2023; Li et al., 2024; Sun et al., 2023). For example, to train a learning-based configuration tuning system, it is necessary to observe how different SQL queries interact with various configuration settings, which requires a large number of diverse SQL workloads. Besides, since SQL can be easily translated into natural language, it also facilitates the construction of datasets for tasks like Text-to-SQL.
@@
High generation cost. Existing methods based on large language models (LLMs) often depend on either proprietary APIs or heavyweight open-source models to achieve reliable quality, leading to high costs in API usage or deployment. For instance, OmniSQL (Li et al., 2025a) demonstrates the feasibility of large-scale SQL generation by using a mixture of locally deployed models, including those with over 30 billion parameters. This approach significantly increases deployment costs.
To address these limitations, we introduce SQL-Factory, a multi-agent framework designed for high-quality and large-scale SQL query generation. SQL-Factory leverages three collaborative agent teams: the Management Team, the Generation Team, and the Expansion Team, each responsible for a distinct aspect of the generation process. The Generation Team explores structurally diverse and expressive SQL patterns using a powerful large-scale language model, ensuring sufficient variety across query forms and structures. The Expansion Team complements this by scaling the workload through a lightweight and cost-effective local model, which generate semantically coherent variants based on representative seed queries, significantly reducing reliance on expensive models. Finally, the Management Team evaluates query quality and schema-level distribution in real time, guiding scheduling decisions to ensure that structurally complex or underrepresented schema components are adequately covered. This division of responsibilities enables SQL-Factory to achieve a promising trade-off between query diversity, generation efficiency, and schema-level balance.
@@ 2.1. Data-Driven Intelligent Database
Another important line of work focuses on converting natural language to SQL (NL2SQL). A wide range of methods (Pourreza and Rafiei, 2023; Wang et al., 2025b; Gao et al., 2024; Zhang et al., 2024; Xu et al., 2025; Dong et al., 2023) train neural models on aligned natural language and SQL query pairs, enabling users to express complex information needs in plain language and receive executable queries as output.
@@ 2.3. LLM-based Multi-Agent System
In the context of databases, LLM-based multi-agent approaches have recently emerged as a powerful solution for tasks requiring high modularity and semantic precision. For instance, in Text-to-SQL translation, multiple agents can collaboratively handle subtasks such as schema linking, query generation, and final SQL refinement, leading to more robust and accurate SQL outputs (Wang et al., 2025b; Xie et al., 2024). In the realm of database anomaly diagnosis, systems like D-Bot (Zhou et al., 2024) leverage several large language models to automatically analyze and identify root causes of anomalies by extracting knowledge from diagnostic documents. Additionally, ROMAS (Huang et al., 2024) introduces a role-based multi-agent framework for database monitoring and planning, featuring agents assigned to specific roles (planner, monitor, and worker) collaboratively facilitate dynamic task allocation, self-monitoring, and adaptive re-planning in complex data environments. These examples collectively demonstrate the potential of multi-agent architectures to enhance reasoning modularity, improve decision accuracy, and integrate system-level feedback more effectively. Despite these advances, the application of LLM-based multi-agent systems to large-scale SQL generation remains underexplored.
@@ 3. System Overview
Generation Team. The Generation Team is designed to maximize structural diversity of generated queries, tackling the problem of repetitive or overly templated query outputs. It comprises the Table Selection Agent and the Generation Agent. The Table Selection Agent realizes schema-aware planning by prioritizing structurally complex tables or those with few synthesized queries. Then, the Generation Agent employs a powerful LLM to generate novel and semantically valid queries for selected tables. These queries often include complex structures such as nested subqueries, aggregations, and multi-table joins. By focusing on pattern exploration, the Generation Team ensures the introduction of high-diversity query structures that guide subsequent expansion.
@@ 4.1.1. Management Agent
Specifically, the system begins in the exploration phase, where the Generation Team synthesizes queries for structurally complex tables or those with few synthesized queries. After a fixed number of generation cycles, control shifts to the Expansion Team to exploit learned patterns. During this exploitation phase, the Management Agent continuously monitors evaluation information from the Critical Agent. If query redundancy increases significantly or executability declines, it triggers a transition back to exploration. This alternating control mechanism allows SQL-Factory to incrementally grow a dataset that remains both diverse and scalable, while preventing the generation process from collapsing into repetitive or overly common query patterns.
@@ 4.1.2. Critical Agent
The Critical Agent serves as the evaluation engine of SQL-Factory. Its primary function is to assess the quality of newly generated or expanded SQL queries on the fly before they are stored into the SQL Pool. This agent operates as a shared component used by both the Generation Team and the Expansion Team, and its evaluation result also plays a crucial role in guiding the Management Agent’s scheduling decisions.
@@
For SQL queries from the Generation Team, the powerful large model ensures a degree of novelty and exploratory coverage. Therefore, we focus primarily on syntactic and semantic validity. Given a set of generated queries $\mathcal{Q}_{t}=\{q_{t}^{1},q_{t}^{2},\dots,q_{t}^{n}\}$ at step $t$, the Critical Agent call the Execution Tool $\text{Tool}_{\text{E}}$ to compile each query on the target database schema and reports a binary executability flag. The overall evaluation for this batch is defined as:
@@
    (5) $\text{Sim}_{\text{Emb}}(q,q^{\prime})=\frac{\text{Emb}(q)\cdot\text{Emb}(q^{\prime})}{||\text{Emb}(q)||\ ||\text{Emb}(q^{\prime})||},$
    where $\text{Emb}(q)$ denotes the vector representation of query $q$ encoded by the BERT encoder, and |||||$\cdot$||| | ⋅ | | represents the norm of the vector.
By integrating the above three similarity metrics, this hybrid scoring function provides a more comprehensive assessment of query similarity. It is particularly effective in identifying subtle redundancies among expanded SQL queries.
Finally, the Critical Agent aggregates the execution validation results from $\text{Tool}_{\text{E}}$ and retrieval results from $\text{Tool}_{\text{R}}$ to produce the final evaluation score.
@@ 4.2.1. Table Selection Agent
The agent operates over two sources of information: (1) the full schema of the target database, including all table names and their respective column types; and (2) a statistical summary of SQL usage derived from the SQL Pool, such as the number of queries for each table. These metrics are extracted and structured via a dedicated Statistical Tool $\text{Tool}_{S}$, which computes table-level query coverage.
@@ 4.2.2. Generation Agent
The Generation Agent is responsible for synthesizing new SQL queries in a schema-aware and open-ended manner. It takes as input the set of tables $\mathcal{T}^{\prime}$ selected by the Table Selection Agent and aims to produce queries that are structurally diverse. This process constitutes the core exploration step in SQL-Factory, enabling the discovery of novel query patterns that cannot be easily obtained by simply modifying existing SQL queries.
@@ 4.3. Expansion Team
The expansion process also follows a two-step workflow. First, the Seed Selection Agent identifies representative queries from the SQL Pool that involve under-covered tables or patterns. Then, the Expansion Agent learns from these queries and generates structurally similar but lexically and semantically varied SQL queries that preserve executability. Since this phase focuses on local exploration around known patterns, it enables SQL-Factory to rapidly increase the number of high-quality queries while maintaining system efficiency. Together, these two agents complements the structural exploration performed by the Generation Team and allows SQL-Factory to produce a large volume of diverse and executable queries at scale with low cost.
@@ 4.3.2. Expansion Agent
The Expansion Agent generates new SQL queries by leveraging representative seed queries retrieved by the Seed Selection Agent. Unlike the Generation Agent, which synthesizes queries from scratch based on schema information, the Expansion Agent focuses on local exploration. Its goal is to generate syntactically diverse and executable queries by learning from the structure or intent of seed queries. This agent enables the system to scale up query volume without incurring the computational cost of large-scale model inference.
@@ 5. Experiments
In this section, we conduct extensive experiments to demonstrate the efficiency of the SQL-Factory framework from multiple perspectives, including detailed analysis on query quality, and its effectiveness on downstream tasks.
@@ 5.1.2. Benchmarks
To evaluate the effectiveness and generality of SQL-Factory, we select four typical database benchmarks for synthesizing queries. (1) TPC-DS (Nambiar and Poess, 2006) is an industry-standard benchmark for decision support systems, featuring a complex schema with 24 tables and various star-join patterns. It serves as a rigorous testbed for generating analytical SQL queries. (2) IMDB (Maas et al., 2011) is a movie-related database, which includes 21 tables on movies and actors. Its real-world structure and moderate size make it suitable for generating realistic join-heavy queries. (3) SPIDER (Yu et al., 2018) is a large-scale text-to-SQL dataset consisting of several databases with highly diverse and normalized schemas. We use the SPIDER schema collection as a target set for multi-domain query generation. (4) BIRD (Li et al., 2023a) is a recent benchmark introduced to evaluate database-instruction reasoning in realistic and complex environments. We use BIRD both as a target for SQL generation and as a downstream testbed for Text-to-SQL evaluation, particularly to assess robustness under schema ambiguity. Across all benchmarks, we parse the schema into structured representations including table names, column types, and foreign key relations. These are used as inputs for SQL-Factory’s Generation and Expansion workflows.
Besides, To evaluate the effectiveness of the generated SQL queries in improving representation learning, we adopt three widely-used query clustering benchmarks: IIT Bombay (Chandra et al., 2015), UB Exam, and PocketData (Kennedy et al., 2016). These datasets contain real-world SQL queries collected from academic and industry environments and their manually annotated cluster labels.
@@ 5.1.3. Baseline
SQL Generation Baselines. To compare the quality and diversity of generated SQL queries, we consider both random generation methods and learning-based generation methods. Specifically, we include SQLSmith (Seltenreich, 2016), a grammar-driven query generator that produces random SQLs over a given schema, and two representative learning-based methods: LearnedSQLGen (Zhang et al., 2022), which learns to generate queries via reinforcement learning, and OmniSQL (Li et al., 2025a), which leverages large language models to synthesize large-scale queries. Additionally, we exclude template-based methods (Nambiar and Poess, 2006; Poess and Floyd, 2000; Leis et al., 2018) from our comparison, as they rely heavily on predefined templates and are not suitable for generating massive and structurally diverse SQL workloads.
@@
Query Clustering Baselines. For the SQL clustering task, we evaluate several established similarity models. These include statistical and heuristic approaches such as Aouiche (Aouiche et al., 2006), Aligon (Aligon et al., 2014), and Makiyama (Makiyama et al., 2015), the pretrained model-based PreQR, and two general-purpose language models: BERT-Base and BERT-Large (Devlin et al., 2019). We further include versions of BERT fine-tuned with contrastive learning on SQL-Factory data to highlight the effectiveness of our generated queries in improving SQL representation learning.
@@ 5.1.4. Implementation Details
Expansion Agent. For the Expansion Agent, we utilize Qwen2.5-Coder-14B (Hui et al., 2024), a locally deployed open-source model optimized for code and SQL-related instruction following capability. It is served using vLLM to support efficient batch decoding at scale, enabling rapid query expansion. It is also use a temperature of 0.8 for diversity.
@@ 5.2.3. Comparative Analysis on Query Diversity
We conduct a comparative study to assess the query diversity achieved by SQL-Factory against three representative SQL generation methods: SQLSmith (Seltenreich, 2016), a grammar-driven random query generator that walks through SQL syntax trees; LearnedSQLGen (Zhang et al., 2022), which employs a reinforcement learning framework guided by cost-based feedback to generate queries through a finite-state abstraction; and OmniSQL (Li et al., 2025a), a LLM-based framework that leverages prompt engineering to generate millions of SQL queries.